# Promoting Constructive Deliberation: Reframing for Receptiveness

### v2

@@ 1 Introduction
NLP research for assisting scalable content moderation has also largely focused on anti-social tasks such as removing toxic speech Pamungkas et al. (2020); Ranasinghe and Zampieri (2021), reducing offensive language with style-transfer Atwell et al. (2022), and removing content that violates community rules Ye et al. (2023). While some studies have begun to pursue content moderation through the lens of promoting pro-social content Ashida and Komachi (2022); Yu et al. (2022); Madaan et al. (2020), promoting constructive deliberation has received relatively little attention. This is a difficult task, involving nuanced and subtle linguistic cues that underlie productive discussions.
Research from communications, psychology and linguistics suggests how we might promote open-mindedness to opposing views. In particular, Yeomans et al. (2020)’s conceptualization of conversational receptiveness focuses on conveying openness while still disagreeing. Others have studied how to measure receptiveness towards disagreeing perspectives Minson et al. (2020); Minson and Chen (2022). However, there has been limited success in translating such theory into practice.
In this work, we seek to bridge NLP research on textual reframing Sharma et al. (2022); Ziems et al. (2022); Madaan et al. (2020); Chakrabarty et al. (2021) with social science theories on receptiveness. Our aim is to promote pro-social deliberation online; particularly to foster productive discussion for contentious issues. We develop a computational framework to signal receptiveness while preserving the original meaning of reframed content, studying the capabilities of LLMs for this task. We illustrate the value of translating social science theory to a computational framework by showing how model generations created with our framework better align with human notions of receptiveness, allowing for more useful content moderation tools.
Models that are more receptive-aligned might allow for content moderation tools that are valuable to users and moderators despite being automated. For moderators, a tool built on our receptiveness framework could allow for more scalable and creative moderation. For users, such a tool might provide learning opportunities, suggesting possible revisions if their post is flagged as unreceptive. Users could learn how to be more receptive from suggestions whether or not they accept them.
@@
We compare reframes generated by our framework to original replies and baselines. We find they are significantly more aligned with what humans find receptive, eliciting relatively low negative emotional reactions, high curiosity to engage with opposing views, and reduced belief that the user is trying to shut down discussion. We also explore the interaction between toxicity and receptiveness reframing, finding that more toxic content benefits more from reframing. We create and share a corpus of receptive-reframed replies, baseline reframes, and human annotations of receptiveness (16.6k reframes and over 9.2k total annotations) [^1] Contact authors for dataset if interested.
@@ 2 Related Work
2.1 Receptiveness in political discussions
@@
We draw on Yeomans et al. (2020)’s concept of conversational receptiveness as it more directly focuses on receptiveness when expressing disagreement: the “extent to which participants in disagreement communicate their willingness to engage with each other’s views.” Their study identified numerous linguistic markers of receptiveness during disagreement, such as hedging and acknowledgement.
@@
2.2 Reframing
@@ 3.1 Reframing strategies for receptiveness
are words or phrases that add ambiguity in a way that softens the force of a statement, thus potentially saving negative face of the receiver. They signal possibility in arguments, allowing writers to present “their claims with caution, and enter into dialogue with their audiences” Hyland (1996). Although hedges may weaken the claim, they tend to strengthen the argument as hedged claims are harder to falsify Meyer (1997). Prior work identifies hedges as a potential signal for receptiveness Minson and Chen (2022); Yeomans et al. (2020). Examples: "may", "perhaps" and "I think".
Acknowledgements
convey understanding and validate another’s feelings, appealing to their positive face. Such affirming speech acts signify to the receiver that the speaker is engaging in good faith, likely resulting in reciprocation Rajadesingan et al. (2021). Examples: "I understand what you mean", "I see where you’re coming from".
@@ Acknowledgements
includes asking for further details, clarifying understanding, and repeating what was heard. Seeking more details may be viewed as delaying expressing disagreement, potentially softening threats to negative face Pomerantz (2021). Further asking for elaboration signals expressed interest in the speaker’s view which is known to increase perceptions of receptiveness Chen et al. (2010). Examples: "what I heard you say was that…", "what did you mean by…".
@@
is often an essential precondition for conflict resolution van Bijnen (2020). Highlighting and making explicit shared beliefs, assumptions and knowledge is known to increase conversational receptiveness in civic-minded workshops to reduce polarization Oliver-Blackburn and Chatham-Carpenter (2023); ListenFirst (2018). While acknowledgement simply conveys understanding of the speaker’s point of view, grounding emphasizes common ground and commonalities between the discussants’ viewpoints. Examples: "we both believe…", "we both think…".
@@
often yields more positive perceptions and increased trustworthiness of the speaker Percival and Pulford (2020); Dunn and Schweitzer (2005). Examples: "thank you for your comment", "I appreciate you bringing this up".
Agreement
typically takes the ‘yes,.. but’ form Pomerantz (2021). This mitigation strategy allows individuals to delay and minimize the force of disagreement with an explicit but nominal token of agreement. This strategy can be viewed as a narrow subset of the grounding strategy. While agreement simply makes explicit some agreement about the speaker’s statements, grounding also includes highlighting even broader shared beliefs (for example, “I am also a libertarian but I disagree on this …”). We separate out this strategy from grounding as our expectation is that highlighting specific agreement with the speaker’s statements may signal more receptiveness than emphasizing broader common ground. Examples: "I agree about …but", "I also think …yet".
@@ 3.2 Task formulation and data
Task. To ground our study, we investigate automated reframing of Reddit discussions. These discussions begin with a post in which someone typically shares a link to news or other content. Next, users comment on the post, after which other users reply to these comments. This initial post-comment-reply plays an important role in establishing the tone and basis for any ensuing discussion, so we focus our reframing efforts here. In particular, could the reply be reframed to convey greater openness to the opposing viewpoint in the comment while preserving its original meaning?
Data. English Reddit data is drawn from the Debagreement dataset Pougué-Biyong et al. (2021) for disagreement detection. It consists of comments and replies from controversial subreddits: r/Brexit, r/democrats, r/Republicans, r/climate, and r/BlackLivesMatters. We exclude r/Brexit due to our US-based annotators being less familiar with Brexit-related people and issues. We select examples that are labeled as disagreement and where the comment and reply are at most 30 words each, since preliminary pilots revealed that longer text increased the cognitive burden on annotators. The average word length of the comments and replies in the Reddit data are 36.15 words and 33.12 words respectively, indicating our length-constrained data is not much smaller than average.
We run all replies through the Perspective API [^2] https://perspectiveapi.com/ and exclude any reply with predicted toxicity above 0.9, as these were typically too toxic to be reframed without significantly altering their original meaning. Further discussion of the interaction between toxicity and reframing is explored in Section 5).
@@ 3.3 Generating reframes
Baseline reframes. We consider two zero-shot prompts with gpt-4 as baselines. For the paraphrase baseline, we prompt to simply paraphrase the reply for each comment-reply pair. This baseline serves to distinguish whether observed differences stem from our specific reframing strategies or any language model generation. For the receptiveness baseline, we directly prompt for receptiveness, without reference to any of our six strategies. This baseline allows us to evaluate our social science-informed reframing framework against a generic prompt to reframe for receptiveness.
In total, we generate six reframes (plus two baseline reframes) for each of $\sim$2k comment-reply pairs in the Debagreement dataset, yielding 16.6k generations. Appendix B lists full prompts and few-shot examples.
@@ 4 Validation Experiments
We next present analyses conducted to validate our reframe generation methods. We assess whether our reframes satisfy three key properties: 1) demonstrate the six distinct strategies for receptiveness; 2) preserve the meaning of original replies being reframed; and 3) remain responsive to the original comment.
4.1 Do reframes reflect the six strategies?
We inspect differences in the most common trigrams found in generated reframes vs. the original replies. We study two probability distributions: (a) $P(t|s)$, the probability of trigram $t$ being generated by strategy $s$, and (b) $P(s|t)$, given observed trigram $t$, the probability of it having been generated by strategy $s$. The former tells us whether the model is adding appropriate linguistic phrases for each strategy, while the latter measures how distinct the generations are for each strategy.
The top trigrams and probabilities are listed in Table 1. For each strategy, trigrams are shown ordered by $P(t|s)$. Regarding (a), we see that the most common trigrams being added do appear to be consistent with each strategy. Further, except for the “thank you for” trigram, trigrams appear less than 10% of the time as part of each strategy, implying that there is reasonable variety in the way the strategy is expressed. In regard to (b), we find that the majority of the strategies are unique. However, the top trigrams in agreement overlap with grounding and acknowledgement. Recall that agreement represents a narrower subset of grounding; while agreement simply makes explicit any agreement with the speaker’s statements, grounding also includes highlighting broad shared beliefs. Further, acknowledgement conveys understanding of the speaker’s statement without necessarily agreeing with them. While the overlap of the “i agree that” trigram with grounding is expected, overlaps with acknowledgement are rather surprising and suggest that better prompts for agreement were needed. We discuss more about this issue in the limitations section. The amount of trigram overlap between the different strategies is shown in Appendix Figure 4.
@@ 4.2 Do the reframes preserve meaning?
Table 2 shows that our reframes exceed both baselines in both form and semantic similarity with the original reply + context. When we compare our reframes to the original reply without the context, both baselines’ form and semantic similarity scores are higher than our reframes. This indicates that our strategies incorporate context, which has been shown to be greatly preferred (Yerukola et al., 2023). While the paraphrase baseline exhibits the lowest ratio of contradictions, all strategies are very low (< 3%); generated replies rarely contradict the original, as desired. We note that these methods do not consider potential added hallucinated content; we are only checking if existing content is retained.
@@ 4.3 Are the reframes contextually relevant?
Table 3: Mean reasonability scores for each strategy, with 95% CI in parentheses. Table 9 shows paired t-test for differences between each strategy and the original.
@@
Surprisingly, the original, human-written replies are perceived to be significantly less reasonable, likely because some of the original replies do not directly respond to the comment (common in social media), while the generated replies almost always take the context of the original comment into account. See Appendix Table 7 for examples. Another possible reason is that annotators may have interpreted reasonable response to mean more receptive which may have affected their ratings.
@@ 5.1 Annotation task
As discussed in Section 3.2, our data consists of comment-reply pairs, along with generated reframes for two baselines and our receptiveness framework. We use 9 versions of each of the 75 original replies: the original, the two baselines, and the six strategy reframes. Annotators are asked to compare the relative receptiveness between two alternative replies to a comment: the original vs. one of our reframes or a baseline generation. Annotation interfaces are shown in Figures 6, 7, and 8 in Appendix G. Annotation is performed in multiple batches with 3-6 workers per example (4.4 on average). Given the subjective nature of this annotation task, inter-annotator agreement is rather low, yielding a Krippendorff’s $\alpha$ score of 0.23. Given this low agreement, we study the 95% confidence interval around the receptiveness scores to evaluate how this disagreement affects variance around the mean receptiveness scores. To evaluate the reframes, we develop a receptiveness index translated from social science literature, as described below.
@@ 5.3 Results
Figure 2: Receptiveness scores for reframes compared to the original replies. The circle indicates the mean. Lines show the 95% CI. Positive values indicate that the reframed reply is more receptive than the original. “B” indicates a baseline.
@@ 5.4 Factor breakdown of receptiveness
We have seen that the framework-generated reframes are more aligned with human perceptions of receptiveness, but why is this the case? We investigate this by examining the receptiveness scores split by the individual factors. To do so, we averaged the values for the two questions in each factor. We can see from Table 4 that while our reframes appear to decrease negative emotions, increase openness, and increase curiosity towards opposing views, they interestingly do not do as well in reducing bias towards the opposing party. We discuss this further in Section 6. Such a factor analysis (drawing upon social science theory) provides deeper insight into the reasoning for the receptiveness scores than would be possible with the direct receptive baseline.

### v3

@@ 1 Introduction
NLP research for assisting scalable content moderation has also largely focused on anti-social tasks such as removing toxic speech Pamungkas et al. (2020); Ranasinghe and Zampieri (2021), reducing offensive language with style-transfer Atwell et al. (2022), and removing content that violates community rules Ye et al. (2023). While some studies have begun to pursue content moderation through the lens of promoting pro-social content Ashida and Komachi (2022); Yu et al. (2022); Madaan et al. (2020); Kim et al. (2022), promoting constructive deliberation has received relatively little attention. This is a difficult task, involving nuanced and subtle linguistic cues that underlie productive discussions.
Research from communications, psychology, and linguistics suggests how we might promote open-mindedness to opposing views. In particular, Yeomans et al. (2020)’s conceptualization of conversational receptiveness focuses on conveying openness while still disagreeing. Others have studied how to measure receptiveness towards disagreeing perspectives Minson et al. (2020); Minson and Chen (2022). However, there has been limited success in translating such theory into practice.
In this work, we seek to bridge NLP research on textual reframing Sharma et al. (2022); Ziems et al. (2022); Madaan et al. (2020); Chakrabarty et al. (2021) with social science theories on receptiveness. Our aim is to promote pro-social deliberation online; particularly to foster productive discussion for contentious issues. We develop a computational framework to signal receptiveness while preserving the original meaning of reframed content, studying the capabilities of large language models (LLMs) for this task. We illustrate the value of translating social science theory to a computational framework by showing how model generations created with our framework better align with human notions of receptiveness, allowing for more useful content moderation tools.
Models that are more receptive-aligned might allow for content moderation tools that are valuable to users and moderators despite being automated. For moderators, a tool built on our receptiveness framework could allow for more scalable and creative moderation. For users, such a tool might provide learning opportunities, suggesting possible revisions if their post is flagged as unreceptive. Users could learn how to be more receptive from suggestions, whether or not they accept them.
@@
We compare reframes generated by our framework to original replies and baselines. We find they are significantly more aligned with what humans find receptive, eliciting relatively low negative emotional reactions, high curiosity to engage with opposing views, and reduced belief that the user is trying to shut down discussion. We also explore the interaction between toxicity and receptiveness reframing, finding that more toxic content benefits more from reframing. We end with broader discussion of receptiveness for content moderation and mitigating polarization. We create and share a corpus of receptive-reframed replies, baseline reframes, and human annotations of receptiveness (16.6k reframes and over 9.2k total annotations) [^1] Data is available at: https://github.com/GauriKambhatla/constructive_deliberation.
@@ 2 Related Work
2.1 Constructive deliberation and receptiveness in communications
@@
We draw on Yeomans et al. (2020)’s conceptulization of conversational receptiveness as it more directly focuses on receptiveness when expressing disagreement: the “extent to which participants in disagreement communicate their willingness to engage with each other’s views.” Their study identified numerous linguistic markers of receptiveness during disagreement, such as hedging and acknowledgement.
@@
2.2 Constructive deliberation and NLP
There has been some NLP research on constructive deliberation. Recent research focuses on grounding, or constructing a shared basis of understanding. Cho and May (2020) create a dataset of two-turn dialogues with a particular grounding speech act (“yes, and”) that implies grounding. They find that a model fine-tuned with this dataset encourages more grounded conversations. Shaikh et al. (2024) study the discrepancies between human grounding and LLM generations with grounding, which they call the “grounding gap”. They come up with a set of grounding acts to quantify LLM attempted grounding, similar to our strategies for receptiveness.
Other work looks at constructive discourse in real-world settings, such as in the space of content moderation. Cho et al. (2024) define metrics for conversational moderation effectiveness, and evaluate language models as moderators. They find that language models that are prompted using insights from social science can provide good feedback to users, but struggle to increase users’ levels of cooperation and respect. Park et al. (2021) study using community norms for better context-specific automated content moderation. They find that explicitly encoding community norms can allow models to detect community norm violations with high performance.
2.3 Reframing in NLP
@@ 3.1 Reframing strategies for receptiveness
are words or phrases that add ambiguity in a way that softens the force of a statement, thus potentially saving negative face of the receiver. They signal possibility in arguments, allowing writers to present “their claims with caution, and enter into dialogue with their audiences” Hyland (1996). Although hedges may weaken the claim, they tend to strengthen the argument as hedged claims are harder to falsify Meyer (1997). Prior work identifies hedges as a potential signal for receptiveness Minson and Chen (2022); Yeomans et al. (2020). Examples of hedges: "may", "perhaps" and "I think".
@@
includes asking for further details, clarifying understanding, and repeating what was heard. Seeking more details may be viewed as delaying expressing disagreement, potentially softening threats to negative face Pomerantz (2021). Further asking for elaboration signals expressed interest in the speaker’s view which is known to increase perceptions of receptiveness Chen et al. (2010). Examples of elaboration: "what I heard you say was that…", "what did you mean by…".
@@
is often an essential precondition for conflict resolution van Bijnen (2020). Highlighting and making explicit shared beliefs, assumptions and knowledge is known to increase conversational receptiveness in civic-minded workshops to reduce polarization Oliver-Blackburn and Chatham-Carpenter (2023); ListenFirst (2018). Acknowledgement and agreement (strategies discussed below) can be viewed as different grounding actions Clark and Schaefer (1989); Traum (1999). We chose to separate these into individual strategies as they may result in different levels or degrees of groundedness which may impact receptiveness. Our expectation is that highlighting specific consensus (agreement strategy) or using affirming speech acts (acknowledgement strategy) may signal receptiveness in a different manner than emphasizing broader common ground. Examples of grounding: "we both believe…", "we both think…".
Acknowledgement
conveys understanding and validate another’s feelings, appealing to their positive face. Such affirming speech acts signify to the receiver that the speaker is engaging in good faith, likely resulting in reciprocation Rajadesingan et al. (2021). While acknowledgement simply conveys understanding of the speaker’s point of view, grounding emphasizes common ground and commonalities between the discussants’ viewpoints. Examples of acknowledgement: "I understand what you mean", "I see where you’re coming from".
Agreement
typically takes the ‘yes,.. but’ form Pomerantz (2021). This mitigation strategy allows individuals to delay and minimize the force of disagreement with an explicit but nominal token of agreement. While agreement simply makes explicit some agreement about the speaker’s statements, grounding also includes highlighting even broader shared beliefs. Examples of agreement: “I am also a libertarian but I disagree on this …”).
@@ Acknowledgement
often yields more positive perceptions and increased trustworthiness of the speaker Percival and Pulford (2020); Dunn and Schweitzer (2005). Examples of gratitude: "thank you for your comment", "I appreciate you bringing this up".
@@ 3.2 Task formulation and data
Task.
To ground our study, we investigate automated reframing of Reddit discussions. These discussions begin with a post in which someone typically shares a link to news or other content. Next, users comment on the post, after which other users reply to these comments. This initial comment-reply plays an important role in establishing the tone and basis for any ensuing discussion, so we focus our reframing efforts here. In particular, could the reply be reframed to convey greater openness to the opposing viewpoint in the comment while preserving its original meaning?
Data.
English Reddit data is drawn from the Debagreement dataset Pougué-Biyong et al. (2021) for disagreement detection. It consists of comments and replies from controversial subreddits: r/Brexit, r/democrats, r/Republicans, r/climate, and r/BlackLivesMatters. We exclude r/Brexit due to our US-based annotators being less familiar with Brexit-related people and issues. We select examples that are labeled as disagreement and where the comment and reply are at most 30 words each, since preliminary pilots revealed that longer text increased the cognitive burden on annotators. The average word length of the comments and replies in the Reddit data are 36.15 words and 33.12 words respectively, indicating our length-constrained data is not much smaller than average.
We run all replies through the Perspective API [^2] https://perspectiveapi.com/ and exclude any reply with predicted toxicity above 0.9, as these were typically too toxic to be reframed without significantly altering their original meaning. Examples of replies for varying toxicity levels is shown in Appendix Table 8. Further discussion of the interaction between toxicity and reframing is explored in Section 5).
@@ 3.3 Generating reframes
Choosing in-context examples.
We choose one comment-reply pair from each subreddit, and reframed responses were written by one of the authors. The final reframes are the result of multiple iterations with feedback from the other authors. In this work, we selected and collectively reviewed real comments from each subreddit for use as our in-context examples. Future work might employ a more calculated strategy to choose examples that better delineate the different strategies.
Baseline reframes.
We consider two zero-shot prompts with gpt-4 as baselines. For the paraphrase baseline, we prompt the model to simply paraphrase the reply for each comment-reply pair. This baseline serves to distinguish whether observed differences stem from our specific reframing strategies or any language model generation. For the receptiveness baseline, we directly prompt for receptiveness, without reference to any of our six strategies. This baseline allows us to evaluate our social science-informed reframing framework against a generic prompt to reframe for receptiveness.
In total, we generate six reframes (plus two baseline reframes) for each of $\sim$2k comment-reply pairs in the subset of the Debagreement dataset that is within our constraints (within the four subreddits, under the 30-word maximum, with a toxicity score under 0.9, and where the comment and reply disagree), yielding 16.6k generations. Appendix B lists full prompts and few-shot examples.
@@ 4 Validation Experiments
We next present analyses conducted to validate our reframe generation methods. We assess whether our reframes satisfy three key properties: 1) demonstrate the six distinct strategies for receptiveness; 2) preserve the meaning of original replies being reframed; and 3) are contextually relevant to the original comment.
@@
4.1 Do reframes reflect the six strategies?
We inspect differences in the most common trigrams found in generated reframes vs. the original replies. We study two probability distributions: (a) $P(t|s)$, the probability of a trigram $t$ being generated by a given strategy $s$, and (b) $P(s|t)$, the probability of given observed trigram $t$, having been generated by a strategy $s$. The former tells us whether the model is adding appropriate linguistic phrases for each strategy, while the latter measures how distinct the generations are for each strategy.
The top trigrams and probabilities are listed in Table 1. For each strategy, trigrams are shown ordered by $P(t|s)$. Regarding (a), we see that the most common trigrams being added do appear to be consistent with each strategy. Further, except for the “thank you for” trigram, trigrams appear less than 10% of the time as part of each strategy, implying that there is reasonable variety in the way the strategy is expressed. In regard to (b), we find that the majority of the strategies are unique. However, the top trigrams in agreement overlap with grounding and acknowledgement. Recall that agreement represents a narrower subset of grounding; while agreement simply makes explicit any agreement with the speaker’s statements, grounding also includes highlighting broad shared beliefs. Further, acknowledgement conveys understanding of the speaker’s statement without necessarily agreeing with them, and is often considerd an aspect of grounding. As a result, the overlap between the three strategies is not surprising, though suggests that our expectations for differentiation (discussed in Section 3.1) did not bear out. This merits further investigation into alternative formulations of these strategies as well as an exploration of others. We discuss more about this issue in Limitations. The amount of trigram overlap between the different strategies is shown in Appendix Figure 4.
@@ 4.2 Do the reframes preserve meaning?
Table 2 shows that our reframes exceed both baselines in both form and semantic similarity with the original reply concatenated with context. When we compare our reframes to the original reply without the context, both baselines’ form and semantic similarity scores are higher than our reframes. This indicates that our strategies incorporate context, which has been shown to be greatly preferred (Yerukola et al., 2023). While the paraphrase baseline exhibits the lowest ratio of contradictions, all strategies are very low (< 3%); generated replies rarely contradict the original, as desired. We note that these methods do not consider potential added hallucinated content; we are only checking if existing content is retained.
@@ 4.3 Are the reframes contextually relevant?
Table 3: Mean reasonability scores for each strategy, with 95% CI. Appendix Table 10 shows paired t-test for differences between each strategy and the original.
@@
Surprisingly, the original, human-written replies are perceived to be significantly less reasonable, likely because some of the original replies do not directly respond to the comment (which is common in social media), while the generated replies almost always take the context of the original comment into account. See Appendix Table 7 for examples. Another possible reason is that annotators may have interpreted reasonable response to mean something more akin to receptive response. We use “reasonability” as a derivative of “plausibility”, which was used in prior work Zhou et al. (2023) for a similar generation validation experiment. We found during pilot studies that “reasonability” was more intuitive than “plausibility” to annotators.
@@ 5.1 Annotation task
As discussed in Section 3.2, our data consists of comment-reply pairs, along with generated reframes for two baselines and our receptiveness strategies. We use 9 versions of each of the 75 original replies: the original, the two baselines, and the six receptive reframes. Annotators are asked to compare the relative receptiveness between two alternative replies to a comment: the original vs. one of our reframes or a baseline generation. Annotation interfaces are shown in Figures 6, 7, and 8 in Appendix G. Annotation is performed in multiple batches with 3-6 workers per example (4.4 on average). We use Krippendorff’s $\alpha$ for interannotator agreement, and get a score of 0.54, which makes sense given the subjective nature of this annotation task. We study the 95% confidence interval around the receptiveness scores to evaluate how this disagreement affects variance around the mean receptiveness scores. To evaluate the reframes, we develop a receptiveness index translated from social science literature, as described below.
@@ 5.3 Results
Figure 2: Receptiveness scores for reframes compared to the original replies. The dot indicates the mean. Lines show the 95% CI. Positive values indicate that the reframed reply is more receptive than the original. “B” indicates a baseline.
@@ 5.4 Factor breakdown of receptiveness
We have seen that our framework-generated reframes are more aligned with human perceptions of receptiveness, but why is this the case? We investigate this by examining the receptiveness scores split by the individual factors. To do so, we averaged the values for the two questions in each factor. We can see from Table 4 that while our reframes appear to decrease negative emotions, increase openness, and increase curiosity towards opposing views, they interestingly do not do as well in reducing bias towards the opposing party. We discuss this further in Section 6. Such a factor analysis (drawing upon social science theory) provides deeper insight into the reasoning for the receptiveness scores than would be possible with the direct receptive baseline.